Equivalent Units and Product Cost Report—FIFO Method

In its first month of operations (May of 2019), Allred Company’s Department 1 incurred charges of $72,000 for direct material (9,000 units), $38,700 for direct labor, and $13,500 for manufacturing overhead. At month-end, 8,500 units had been finished and transferred out. Those units remaining were finished with respect to material but only 40% finished with respect to conversion.

Assuming Allred uses the FIFO method and that materials are added at the beginning of the process and conversion costs are incurred evenly throughout the period, compute the following:

  1. The equivalent units for materials and conversion costs.

  2. The cost per equivalent unit for materials and conversion costs.

  3. The total cost assigned to the units transferred out.

  4. The total cost assigned to the ending inventory.

  5. Prove that your solutions to requirements (c) and (d) sum to the total costs to be accounted for.

Allred Company, Department 1
Equivalent Units Calculations, May 2019
Equivalent Units
% in
May
Direct
Materials
% in
May
Conversion
Costs
Complete/Transferred
Beginning inventory Answer Answer Answer Answer Answer
Started & completed Answer Answer Answer Answer Answer
Ending inventory Answer Answer Answer Answer Answer
Total Answer Answer

a.

Answer a.


Product Cost Report
Direct
Materials
Conversion
Costs
Beginning Inventory Answer Answer Answer
Current Answer Answer Answer
Total Costs to Account For Answer Answer Answer
÷ Total Equivalent Units Answer Answer
Average cost / Equivalent unit Answer b. Answer b.
Beginning inventory
Costs incurred in Month 0 Answer
Costs incurred in Month 1 Answer
Started and finished Answer
Cost of Goods Manufactured Answer c.
Ending Inventory:
Direct Materials Answer
Conversion costs Answer
Cost of Ending Inventory Answer d.
Total Costs Allocated Answer e.
Answer #1

Calculation for Ending WIP:-

Beginning WIP + Started = Ending WIP + Completed & Transferred

Since this is the 1st month of production there is no beginning WIP.

0 + 9,000 = Ending WIP + 8,500

Ending WIP = 9,000 - 8,500

= 500

Equivalent Units for Direct Materials :-

Completed & Transferred Out 8,500

Ending WIP (100% Added) 500

Equivalent units 9,500 units

Equivalent units for Conversion Cost:-

Completed & Transferred Out 8,500

Ending WIP (40% Completed) (500*40%) 200

Equivalent units 8,700 units

Cost per Equivalent units for Materials:-

Material cost/Equivalent units for Materials

= $72,000/9000

= $8/Unit

Cost per Equivalent units for Conversion Cost:-

Conversion cost/Equivalent units for conversion

($38,700+$13,500)/8700

= $6/Units

Cost assigned to Units Transferred Out:-

Direct Materials = $8*8,500 = $68,000

Conversion cost = $6*8,500 = $51,000

Total cost assigned = $119,000

Cost Assigned to Ending WIP:-

Direct Materials = $8*500 = $4,000

Conversion Cost = $6*200 = $1,200

Total cost assigned = $5,200

Total cost to be Accounted for:-

Direct Materials = $72,000

Direct Labor = $38,700

Overhead = $13,500

Total cost = $124,200

Cost Assigned to Transferred out = $119,000

Cost Assigned to Ending WIP = $ 5,200

Total Cost Assigned = $124,200